OpenStudy (anonymous):

hey..... A state patrol officer saw a car start from rest at a highway on-ramp. She radioed ahead to another officer 30 mi along the highway. When the car reached the location of the second officer 26 min later, it was clocked going 60 mi/hr. The driver of the car was given a ticket for exceeding the 60/mi/hr speed limit. Why can the officer conclude that the drive exceeded the speed limit?

OpenStudy (tkhunny):

HINT: 60 mph = 1 mile/minute

OpenStudy (anonymous):

Can you please explain it? I really dont get it

OpenStudy (lukebluefive):

If the car was going 60mph, or 1 mile a minute, how long would it have taken the car to travel 30 miles?

OpenStudy (anonymous):

30min

OpenStudy (tkhunny):

...and...

OpenStudy (anonymous):

so 30 miles in 30 min?

OpenStudy (lukebluefive):

Right, so how is the only way the car could've travelled 30 miles in less than 30 minutes?

OpenStudy (anonymous):

if it went over 30mph?

OpenStudy (lukebluefive):

Well, a mile a minute is 60 mph, so 30 miles would take 30 minutes to travel, going that speed. So the car wasn't just going over 30 mph, it was going over...?

OpenStudy (anonymous):

60mph

OpenStudy (lukebluefive):

Exactly, which is what you wanted to show.

OpenStudy (anonymous):

But it is telling me to use eithe Rolle's theorem, mean value theorem or extreme value theorem which one do I know to use?

OpenStudy (anonymous):

and it wants to know at some time c, the car was traveling at ___mph

OpenStudy (lukebluefive):

I think you'd use Rolle's Theorem to find that out.

OpenStudy (anonymous):

can you help me walk me through it?

OpenStudy (lukebluefive):

Actually, it's the Mean Value Theorem... Rolle's Theorem is just a specific type of it, but w/e. Let me see if I can get brushed up enough to help you.

OpenStudy (anonymous):

thank you

OpenStudy (tkhunny):

Simplified - The average speed must have been accomplished at some time during the measurement. Other Words - If it's continuous, it can't be the average if you don't manage it at least once.

OpenStudy (lukebluefive):

Well, first, I'd find the average speed using the following equation: \[speed = \frac{distance} {time}\]

OpenStudy (anonymous):

ok

OpenStudy (lukebluefive):

Since the car travelled 30 miles in 26 minutes, this would be: \[speed = \frac{30 miles}{26 minutes} = \frac{15miles}{13minutes} \times \frac{60minutes}{1hour} = \frac{900miles}{13hours} \approx 69 mph\]

OpenStudy (anonymous):

thank youuu

OpenStudy (lukebluefive):

You're welcome. Let me know if you need help in the future.